OpenStudy (anonymous):

solve -8x+y=46, 9x+17y=57

OpenStudy (anonymous):

\[-8x+y=46\]\[ 9x+17y=57\]this can easily be solved using the substitution method, solve for y in the first equation then plug it into the y in the second equation; next just plug in the resulting x value in to either equations and find y :) heres how: \[-8x+y=46~~\implies~~y=8x+46\]\[9x+17y=57~~\implies~~9x+17(8x+46)=57\]\[\text{simplify and find the x --->}\]\[9x+17(8x+46)=57\implies~~9x+136x+782=57~~\implies~~145x+782=57\]\[\implies~~145x=-725~~\implies~~x=-5\]\[\text{plug in the x into either equation --->}\]\[-8(-5)+y=46~~\implies~~40+y=46~~\implies~~y=6\] \[\text{so now your solutions are x=-5 and y=6 :)}\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you're welcome, anytime! :)